OpenStudy (haileyd):

Abandoned mines frequently fill with water. Before an abandoned mine can be reopened, the water must be pumped out. The size of pump required depends on the depth of the mine. If pumping out a mine that is D feet deep requires a pump that pumps a minimum of D^2/25+ 4D – 250 gallons per minute, pumping out a mine that is 150 feet deep would require a pump that pumps a minimum of how many gallons per minute?

OpenStudy (doodleduck8):

All you have to do here is plug in 150 for D. (150)^2/25 + 4(150) - 250 = answer 1250

OpenStudy (doodleduck8):

The correct answer is 1,250.If you substitute D with 150 in the expression, you get + 4(150) - 250 = + 600 - 250 = 1,250.